OpenStudy (vanillaangelic):

The length of a rectangle is 12 inches, and its diagonal measures 13 inches. Find the area of that rectangle. Please explain step by step.

OpenStudy (anonymous):

\[\sqrt{13^2-12^2}=5 \] 12*5 = 60 sq in

OpenStudy (vanillaangelic):

Thank you.